OpenStudy (anonymous):

Solve the simultaneous equations using the elimination method: 4x-3y=5 2x-3y=3 Help PLEASE!

OpenStudy (anonymous):

3 appears in both equations. The 3's both have the same sign --> minus. Use the rule same signs subtract, different signs add. Subtract top from bottom to get: 2x - 0y = 2 2x = 2 x = 1 Have a nice day. :)

OpenStudy (anonymous):

Ops the next one's 2x+3y=3 :\

OpenStudy (anonymous):

Ok. Sorry I took this long... 3 still appears in both equations but now they have different signs. The 3 can be eliminated by using the rule: different signs add. So adding the top and bottom equations we get: 6x + 0 = 8 6x = 8 x = 8/6 x = 4/3 Now to get y we sub x = 4/3 into one of the original equations: Let's do that with eqn. 2: 2x + 3y = 3 2(4/3) + 3y = 3 8/3 + 3y = 3 --> Multiply everything by 3 to get rid of the fraction: 8 + 9y = 9 9y = 1 y = 1/9 There we go. :)